Imagine that you are sampling a bean population composed of red (R) and white (W) individuals. You find the following number of offspring from each of the three genotypes: RR=12, RW=60, WW=24. Find the relative fitness (w) and calculate the selection coefficient (s) for each genotype. What type of selection does this represent? What should be the eventual frequencies of the 2 alleles R and W, given your fitness estimates? Do you get the same result when you do the computer simulation?
Answer #1

Relative Fitness (w) is the ratio of survival or reproductive rate of a genotype (or phenotype) to maximum survival or reproductive rate of another genotypes in the population.

The genotype RW is more fit.

So, relative fitness (w),

w_{RR}=\frac{12}{60}=0.2

WRW = SIS

w_{WW}=\frac{24}{60}=0.4

Selection coefficient (s),

SRR = 1- WRR = 1-0.2 = 0.8

SRW = 1- WRW =1-1=0

Sww = 1- www = 1-0.4 = 0.6

This represents stabilizing selection.

RR 12 + RW 60 ww. 24 + - 96 20+ 12 - 42) (20+12 = 54) 0 fregency of Rallel - 42 -0.4 frequency of het allel : 54 . 0.6 96

We were unable to transcribe this image

WRW = SIS

We were unable to transcribe this image

SRR = 1- WRR = 1-0.2 = 0.8

SRW = 1- WRW =1-1=0

Sww = 1- www = 1-0.4 = 0.6

RR 12 + RW 60 ww. 24 + - 96 20+ 12 - 42) (20+12 = 54) 0 fregency of Rallel - 42 -0.4 frequency of het allel : 54 . 0.6 96